### v1

# SKA-Bench: A Fine-Grained Benchmark for Evaluating Structured Knowledge Understanding of LLMs

Abstract
Although large language models (LLMs) have made significant progress in understanding Structured Knowledge (SK) like KG and Table, existing evaluations for SK understanding are non-rigorous (i.e., lacking evaluations of specific capabilities) and focus on a single type of SK. Therefore, we aim to propose a more comprehensive and rigorous structured knowledge understanding benchmark to diagnose the shortcomings of LLMs. In this paper, we introduce SKA-Bench, a Structured Knowledge Augmented QA Benchmark that encompasses four widely used structured knowledge forms: KG, Table, KG+Text, and Table+Text. We utilize a three-stage pipeline to construct SKA-Bench instances, which includes a question, an answer, positive knowledge units, and noisy knowledge units. To evaluate the SK understanding capabilities of LLMs in a fine-grained manner, we expand the instances into four fundamental ability testbeds: Noise Robustness, Order Insensitivity, Information Integration, and Negative Rejection. Empirical evaluations on 8 representative LLMs, including the advanced DeepSeek-R1, indicate that existing LLMs still face significant challenges in understanding structured knowledge, and their performance is influenced by factors such as the amount of noise, the order of knowledge units, and hallucination phenomenon. Our dataset and code are available at https://github.com/Lza12a/SKA-Bench.
SKA-Bench: A Fine-Grained Benchmark for Evaluating Structured Knowledge Understanding of LLMs
Zhiqiang Liu1,3, Enpei Niu1,3, Yin Hua1,3, Mengshu Sun4, Lei Liang4, Huajun Chen2,3, Wen Zhang1,3thanks: Corresponding Author. 1School of Software Technology, Zhejiang University 2College of Computer Science and Technology, Zhejiang University 3ZJU-Ant Group Joint Lab of Knowledge Graph 4Ant Group {zhiqiangliu,zhang.wen}@zju.edu.cn
1 Introduction
With the rapid development of large language mo-
dels (LLMs) OpenAI (2023); Dubey et al. (2024), Structured Knowledge (SK), such as knowledge graphs (KG) Bollacker et al. (2008) and tables, still remain essential due to their systematic and rigorous organizational formats. On the one hand, structured knowledge is usually present in various real-world scenarios (e.g., financial reports with numerous tables Chen et al. (2021) and product knowledge graphs Wu et al. (2024)), thus serving as a significant knowledge base for existing LLM systems Liang et al. (2024); Wang et al. (2025). On the other hand, due to their well-organized structure and intensive knowledge characteristics, structured knowledge is also widely utilized to improve the inference-time performances of LLMs Li et al. (2024a, b); Guan et al. (2024). Consequently, evaluating the ability of LLMs to understand structured knowledge is a crucial research topic.
Figure 1: The components of a SKA-Bench instance and how to further construct the four ability testbeds for evaluating structured knowledge understanding.
Unlike common unstructured text understanding tasks Guo et al. (2023), LLMs still face significant challenges Fang et al. (2024) in understanding structured knowledge. This is because LLMs need to capture long-distance contextual dependencies as well as complex relationships and hierarchical structures from the given structured knowledge. However, existing benchmarks Pasupat and Liang (2015); Wu et al. (2025); Talmor and Berant (2018); Wu et al. (2024) for evaluating structured knowledge understanding suffer from limitations, including the lack of detailed reasoning path annotations or sufficiently long structured knowledge bases, making it difficult to thoroughly diagnose the shortcomings of LLMs in structured knowledge understanding. Moreover, these datasets primarily focus on single data types, including tables Pasupat and Liang (2015); Wu et al. (2025), knowledge graphs Talmor and Berant (2018), or hybrid Chen et al. (2020b); Wu et al. (2024) formats, which restrict their coverage and fail to fully reflect the comprehensive understanding abilities of the models. Therefore, there is an urgent need for a diverse and fine-grained dataset to comprehensively evaluate LLMs and identify potential bottlenecks in their structured knowledge understanding capabilities.
To this end, we construct a fine-grained Stru-ctured Knowledge Augmented QA Benchmark, SKA-Bench, which consists of 921 SKA-QA instances and covers four widely used types of structured data. To ensure the quality and complexity of the instances, we propose a novel three-stage construct pipeline for precise positive knowledge unit annotation and the synthesis of long structured knowledge. As illustrated in Fig. 1, SKA-Bench instances are composed of a question, an answer, positive knowledge units, and noisy knowledge units, which endow SKA-Bench with strong scalability. Ultimately, based on the different compositions of SK units as the given structured knowledge bases, we expand these instances into four distinct testbeds, each targeting a fundamental capability required for understanding SK: Noise Robustness, Order Insensitivity, Information Integration, and Negative Rejection for comprehensively diagnosing the shortcomings of LLMs in SK understanding.
We conduct empirical evaluations on 8 representative LLMs. Even advanced LLMs like DeepSeek-R1 continue to face challenges in SK understanding, with their performance significantly influenced by the amount of noise and the order of knowledge units. Moreover, its negative rejection ability is even weaker than that of certain LLMs with 7B parameters. We hope that SKA-Bench can serve as a comprehensive and rigorous benchmark to accelerate the progress of LLMs in understanding and reasoning over structured knowledge.
2 Related Work
Evaluation for Structured Knowledge Understanding.
Current structured knowledge understanding evaluations often focus on knowledge graphs Yih et al. (2016); Talmor and Berant (2018); He et al. (2024) and tables Pasupat and Liang (2015); Zhong et al. (2017); Wu et al. (2025). Earlier Table QA datasets, such as WTQ Pasupat and Liang (2015), WikiSQL Zhong et al. (2017), and TabFact Chen et al. (2020a) require to retrieve several specific table cells with less than 3 hops, posing limited challenges for LLMs. Recently, Wu et al. (2025) propose a more complex Table QA benchmark TableBench for LLM evaluation. However, we believe that the existing evaluations aren’t comprehensive enough. On the one hand, the tables in these Table QA datasets are relatively short (average <16.7 rows), making it difficult to evaluate the ability of LLMs to handle long structured knowledge. On the other hand, these datasets lack detailed reasoning path annotations, limiting their utility in fine-grained evaluation of LLMs’ understanding capabilities. For existing KGQA datasets, such as WebQSP Yih et al. (2016), CWQ Talmor and Berant (2018), and GraphQA He et al. (2024), they are constructed upon large-scale KGs, thus providing a foundation for creating long and complex KG understanding datasets. But they also lack precise positive triple annotations for systematic evaluation and analysis.
Figure 2: The construct pipeline to generate SKA-Bench instance, which consists of structured knowledge augmented question & answer (SKA-QA), positive knowledge units and noisy structured units.
Evaluation for Semi-structured Knowledge Understanding.
To more effectively evaluate the understanding of heterogeneous data, the research community has begun to focus on semi-structured knowledge Chen et al. (2020b); Zhu et al. (2021); Wu et al. (2024) (i.e., structured data integrated with unstructured textual documents). The semi-structured dataset HybridQA Chen et al. (2020b), which combines table and textual data, was first proposed. Subsequently, TAT-QA Zhu et al. (2021) and FinQA Chen et al. (2021) extend the evaluation of understanding and reasoning to more realistic scenarios based on this data format. In addition, STaRK Wu et al. (2024) dataset based on KG and textual knowledge bases introduces a new retrieval and reasoning challenge for LLMs. However, these hybridQA datasets are also limited by relatively short length of tables or lack of precise annotations, making them challenging for systematic evaluation.
Based on the above considerations, we believe that offering a diverse, fine-grained, and complex benchmark is valuable for thoroughly evaluating LLMs’ structured knowledge understanding ability.
3 SKA-Bench
3.1 Problem Definition
To comprehensively evaluate the ability of LLMs in structured knowledge understanding, SKA-Bench incorporates four common types of (semi-)structured data: Knowledge Graph (KG) $\mathcal{G}$, Table $\mathcal{T}$, Knowledge Graph with Textual Documents $\mathcal{G}\cup\mathcal{D}$, and Table with Textual Documents $\mathcal{T}\cup\mathcal{D}$. Following the most existing LLM evaluations Chang et al. (2024); Guo et al. (2023), SKA-Bench also adopts a question-answering (QA) format. For a given question $\mathcal{Q}$ and its corresponding structured knowledge $\mathcal{SK}\in$$\{\mathcal{G},\mathcal{T},\mathcal{G}\cup\mathcal{D},\mathcal{T}\cup\mathcal{D}\}$, the LLM $f_{\theta}$ aims to generate the correct answer $\mathcal{A}$, such that $\mathcal{A}=f_{\theta}(\mathcal{Q},\mathcal{SK})$. We hypothesis that LLMs must accurately understand structured knowledge (SK) as a prerequisite for generating correct answers. Therefore, this task format can thoroughly evaluate the SK understanding capabilities of LLMs.
3.2 SKA-Bench Construction
In this section, we detail the construction process of Structured Knowledge Augmented Benchmark (SKA-Bench), which includes three stages: SKA-QA pairs collection, iterative positive units annotation and noisy units synthesis, shown in Fig 2.
3.2.1 SKA-QA Pairs Collections
Knowledge Graph. We randomly select 900 samples from the test set of KGQA datasets: WebQuestionsSP (WebQSPYih et al. (2016) and ComplexWebQuestions (CWQTalmor and Berant (2018) as the initial SKA-QA pairs of KG subset. These two datasets cover 7 common KG relational patterns Dutt et al. (2023) and are both based on widely used Freebase KG Bollacker et al. (2008). For each QA sample, we extract up to 4-hop subgraph of the topic entities Jiang et al. (2023b) in Freebase as the structured knowledge base.
Table. We randomly select 700 samples from the widely used Table QA dataset WTQ Pasupat and Liang (2015) and TableBench Wu et al. (2025) with multi-domain, multi-hop question as the initial SKA-QA pairs of Table subset. And our selected tables contain at least 6 columns and 8 rows to facilitate the subsequent synthesis of noisy data.
KG with Textual Documents. We choose the STaRK Wu et al. (2024) dataset, which is constructed based on both textual and relational knowledge bases. Specifically, we randomly select 300 QA samples from both STaRK-Prime and STaRK-Amazon. For each QA sample, we extract the 2-hop subgraph of the answer entity and the textual descriptions of neighboring nodes within subgraph as the corresponding structured knowledge base. Additionally, we remove SKA-QA pairs where the number of triples in subgraph is less than 200.
Figure 3: Four Instances from different subsets of SKA-Bench: LLMs need to understand structured knowledge, then select relevant knowledge units to get the answer.
Table with Textual Documents. For this hybrid data, we also require that QA tasks simultaneously utilize multiple data types. Therefore, we select 200 samples from HybridQA Chen et al. (2020b) dataset as a subset. This dataset necessitates reasoning based on heterogeneous knowledge sources and has been widely used in the research community Rogers et al. (2023); Fang et al. (2024).
After obtaining the above four types of SKA-QA pairs, we perform a fine-grained split for structured knowledge. Specifically, we regard the triples $\mathcal{F}$ in the KG $\mathcal{G}$ and the rows $\mathcal{R}$ in the tables $\mathcal{T}$ into individual “structured knowledge units”, represented as $\mathcal{G}=\{\mathcal{F}_{i}\}_{i=1}^{n}$ and $\mathcal{T}=\mathcal{H}\cup\{\mathcal{R}_{j}\}_{j=1}^{n}$. For the table header $\mathcal{H}$, they are separated out independently to preserve the semantic integrity of the table. As for the textual data, we retain the original paragraph-level split in the initial SKA-QA pairs.
3.2.2 Iterative Positive Units Annotation
We invite three human experts with computer science backgrounds to perform positive units annotation. Specifically, we require the human experts to accurately identify the positive units required to derive the answer to the given question. Furthermore, the annotation process need to adhere to the following requirements: (1) if the answer is wrong, delete the sample directly; (2) if the question involves multiple answers, all positive units require to obtain the answers should be annotated; (3) for the Table subset and Table+Text subset, if the question needs to perform numerical analysis on the entire table, the corresponding SKA-QA pairs should either be removed or the question should be modified; (4) if the tables in the Table subset and Table+Text subset are order-dependent (i.e., modifying the row order would result in semantic errors in the table), this sample should be removed; (5) for the KG+Text subset and Table+Text subset, if question only utilizes one type of knowledge source, the question should be modified or removed.
After each round of annotation, we query the LLM (utilizing DeepSeek-v3 DeepSeek-AI et al. (2024)) to determine whether annotated positive units can derive the answer to the given question. If the response is “No”, re-annotation is performed. The iterative annotation process continues until more than 95% of the samples receive a “Yes” response, at which point the iteration is terminated.
Subset #avg $\mathcal{Q}$ token #avg $\mathcal{A}$ num  #num P (SK/unSK)  #avg P token #num N  #data Expert Time
SKA-Bench-KG 15.75 1.96 4.25 16.77 4541.39 233 5.9 min
SKA-Bench-Table 23.31 1.10 3.40 30.88 1521.83 295 3.6 min
SKA-Bench-KG+Text 30.76 1.86 2.53/1.92 22.31/1053.55 417.29/79.84 195 6.8 min
SKA-Bench-Table+Text 22.41 1.01 1.17/1.28 28.37/203.78 1144.55/661.90 198 5.8 min
Table 1: The data statistics of four subsets in SKA-Bench. ‘#num P’ and ‘#num N’ refer to the average number of positive units and noisy units. And ‘#avg P token’ denotes the average number of tokens in positive units. ‘#data’ refers to the numbers of instances in each subsets. The calculation of tokens is based on GPT-4o’s tokenizer. ‘Expert Time’ refers to the median time for each question spent on annotation by human experts.
Figure 4: The distribution of the number of positive units across four SKA-Bench subsets.
3.2.3 Noisy Units Construction
For KG subset and KG+Text subset, we regard all knowledge units in the knowledge base except for the positive units as noisy units. The raw tables in the Table subset and Table+Text subset are typically short (average <17.9 rows), making it hard to comprehensively evaluate the table knowledge understanding of LLMs. Therefore, we introduce an automated noisy data synthesis process as follows.
First, we leverage LLMs with existing SKA-QA instances to generate noisy units. To ensure the diversity of synthesized units, we alternately use GPT-4o OpenAI (2023) and DeepSeek-v3 DeepSeek-AI et al. (2024) during this process. Meanwhile, we also need to ensure that the synthesized noisy units do not affect the correctness of the answers. To achieve this, we prompt LLM (utilizing DeepSeek-v3) with QA and positive units to derive the “conditions” that must be satisfied by the rows for answering the question. LLM then verifies whether the generated noisy units meet these “conditions”. If the response is “Yes”, the noisy units need to be re-generated by LLMs. After the noise synthesis process, three human experts conduct a manual review of Table subset and Table+Text subset to evaluate whether the synthetic noise is unsafe and affect the original answers. The review results show that the accuracy rate is 92.5%, and erroneous noise has been deleted.
3.3 Dataset Statistic
Through the aforementioned construction pipeline, we have completed constructing SKA-Bench instances as shown in Fig. 3, which consist of four main components: question, answer, positive knowledge units, and noisy knowledge units. Detailed statistics are presented in Table 1. Additionally, we detail the human annotation results, i.e., the number of positive units across the four subsets, as shown in the Fig. 4.
3.4 Testbeds Construction
As shown in Fig. 1, inspired by Chen et al. (2024) in text understanding evaluation, we construct the four testbeds based on SKA-Bench instances to evaluate the following fundamental capabilities of LLMs in structured knowledge (SK) understanding:
[…]
Subset #num SK #num unSK #token
SKA-Bench-KG-1k 34.23 - 637.64
SKA-Bench-KG-4k 150.34 - 2831.40
SKA-Bench-KG-12k 604.35 - 11394.18
SKA-Bench-KG-24k 1167.19 - 22036.82
SKA-Bench-Table-1k 29.39 - 777.45
SKA-Bench-Table-4k 130.39 - 3268.45
SKA-Bench-Table-12k 488.00 - 12054.15
SKA-Bench-Table-24k 958.78 - 23595.51
SKA-Bench-KG+Text-4k 11.37 2.91 3172.54
SKA-Bench-KG+Text-12k 40.84 6.79 7417.67
SKA-Bench-KG+Text-24k 153.45 19.11 21644.20
SKA-Bench-Table+Text-4k 25.82 14.58 3510.74
SKA-Bench-Table+Text-12k 75.81 119.01 11899.54
SKA-Bench-Table+Text-24k 165.81 369.01 23070.81
Table 2: The data statistics for subsets with different scales of structured knowledge (SK) bases. ‘#num SK’ represents the number of structured knowledge units, ‘#num unSK’ represents the number of unstructured knowledge units in hybrid subsets. And ‘#token’ represents the total number of tokens in the knowledge bases.
Model KG Table KG+Text Table+Text
    1k     4k    12k    24k     1k     4k    12k    24k     4k    12k    24k     4k    12k    24k
Open Source LLMs
Llama3.1-8B 67.53 58.19 45.86 42.34 27.56 23.52 22.16 13.05 67.02 58.89 49.28 30.27 18.44 12.48
TableGPT-2 78.93 66.76 53.14 48.49 24.40 24.05 20.09 16.02 64.84 55.16 46.92 35.91 25.63 25.60
Qwen2.5-7B 72.45 60.00 47.98 40.97 36.69 32.04 30.45 28.68 76.51 62.82 51.83 38.49 36.00 28.56
GLM4-9B 82.95 66.04 52.75 49.95 19.55 17.71 16.77 17.26 75.39 65.14 55.29 32.13 33.65 30.13
Mistral-7B 59.04 60.34 47.98 45.20 17.67 18.11 16.91 16.19 69.37 66.97 53.54 29.21 25.40 15.83
Advanced General-Purpose LLMs
DeepSeek-v3 85.06 73.93 65.85 59.08 54.42 51.83 47.58 45.57 77.12 74.96 68.87 55.64 53.61 48.55
GPT-4o 85.33 73.42 63.04 58.61 51.39 45.18 40.55 38.24 77.38 73.53 67.39 56.52 53.28 51.97
DeepSeek-R1 89.95 81.58 70.32 64.67 61.96 61.88 61.02 58.24 83.14 78.67 71.92 62.24 57.62 56.97
Table 3: Detailed results of noise robustness analysis. The best results are marked bold and the second-best results are underlined in each column. Cells with darker colors indicate the better performance under this subset.
[…]
Figure 5: Overall noise robustness results on four subsets. ‘Average’ represents the average results across on all results of four subsets.
$\bullet$ Negative Rejection. We hope LLMs should minimize the occurrence of hallucination phenomena Huang et al. (2023) as much as possible when understanding SK. To evaluate this, we construct a negative rejection testbed, where the input SK base consists solely of noisy knowledge units. In this scenario, the LLMs are expected to respond with “I don’t know” or other rejection signals. In this testbed, the provided SK don’t contain any positive units, ensuring broken reasoning paths to evaluate the refusal capability of LLMs. The dataset size and the ordering of knowledge units follow the same settings as “Information Integration” testbed.
Model KG Table KG+Text Table+Text
prefix random suffix original prefix random suffix prefix random suffix original prefix random suffix
Open Source LLMs
Llama3.1-8B 55.07 58.19 65.85 23.52 22.71 19.47 24.57 61.85 58.89 62.55 18.44 22.37 18.53 24.41
TableGPT-2 82.07 66.76 77.36 24.05 26.40 17.25 21.62 57.47 55.16 54.53 25.63 36.75 24.44 28.03
Qwen2.5-7B 78.60 60.00 75.70 32.04 33.26 24.07 31.38 64.89 62.82 67.74 36.00 48.29 29.37 37.46
GLM4-9B 81.30 66.04 82.55 17.71 21.15 12.38 16.22 70.05 65.14 69.34 33.65 41.20 23.89 31.14
Mistral-7B 73.28 60.34 64.30 18.11 21.32 14.76 15.92 63.19 66.97 66.36 25.40 33.16 15.44 27.78
Advanced General-Purpose LLMs
DeepSeek-v3 84.40 73.93 87.52 51.83 49.32 44.75 51.31 76.81 74.96 76.41 53.61 55.80 47.02 49.06
GPT-4o 81.75 73.42 83.69 45.18 45.62 40.47 43.33 74.88 73.53 74.98 53.28 54.88 47.72 52.23
DeepSeek-R1 89.90 81.58 89.40 61.88 67.11 61.63 64.36 79.60 78.67 81.12 57.62 59.28 53.04 57.97
Table 4: Results of order insensitivity analysis. The best results are marked bold and the second-best results are underlined in each column. Cells with darker colors indicate the better performance under this subset.
Figure 6: Correlation coefficients of overall F1 and standard deviation across 4 SK types under noise robustness testbed.
4 Experiments
4.1 Experimental Settings
Models.
Our evaluation is based on popular large language models (LLMs) with a context window of at least 24k tokens. Our evaluated LLMs include advanced general-purpose LLMs: DeepSeek-v3 DeepSeek-AI et al. (2024), GPT-4o OpenAI (2023), DeepSeek-R1 DeepSeek-AI et al. (2025) and common open-source LLMs: Llama-3.1-8B-Instruct Dubey et al. (2024), Qwen2.5-7B-Instruct Team (2024), GLM4-9B-Chat Zeng et al. (2024), and Mistral-7B-Instruct-v0.3 Jiang et al. (2023a). Moreover, we also evaluate the table-specific open-source LLM TableGPT-2 Su et al. (2024), which are trained based on Qwen2.5-7B.
Evaluation Metric.
To evaluate SKA-Bench, we utilize the macro-F1 score as our metrics, which measures the agreement between the predicted answer list and the gold answer list. For the negative rejection testbed, we adopt the “Rejection Rate” as the evaluation metric, which reflects the proportion of instances where the LLMs provide a refusal response out of the total number of test samples when only noisy knowledge units are provided.
Figure 7: Overall order insensitivity results on four subsets. ‘Average’ represents the average results across on all results of four subsets.
4.2 Noise Robustness Analysis
From the results in Table 3, it can be observed that as the length of SK input to LLM increases, the performance degradation across various LLMs becomes significantly pronounced. In particular, Llama3.1-8B exhibits a dramatic decline of up to 58.77% when evaluated on the Table+Text subset from 4k to 24k scale. DeepSeek-R1 demonstrates optimal results across all subsets, whereas GLM4-9B and Qwen2.5-7B achieve relatively competitive performance among the smaller models with the 7-10B parameters.
To further analyze model performance on different data types, we present the mean and standard deviation of F1 scores, and their correlation matrix across 4 subsets, as shown in Fig. 5 and 6. We can observe that the performance trends of different LLMs across 4 SK types are similar in general, with all spearman $\rho>0.64$. However, there are significant differences in the noise robustness of different LLMs across 4 SK types as shown in Fig. 6(B). GLM4-9B can perform well on the KG subset but struggles to understand Table data, and TableGPT-2 leverages large-scale table-related task instruction fine-tuning on the base model Qwen2.5-7B, but its performance on both the Table and Table+Text subsets is less satisfactory. We attribute this to the loss of generalization capabilities due to its specialized training, making it less adaptable to unseen table formats and other data modalities. Furthermore, we observe that DeepSeek-R1 achieves the lowest average standard deviation, exhibiting the strongest noise robustness. This suggests that current LLMs are evolving towards greater robustness against noise.
Figure 8: Correlation coefficients of overall F1 and standard deviation across 4 SK types in order insensitivity testbed.
4.3 Order Insensitivity Analysis
Figure 9: Information Integration results on four subsets demonstrates the variation in F1 score as the number of required positive units increases.
From the results in Table 4, we can observe that when the positive units are concentrated in the prefix or suffix of the structured knowledge base, models tend to focus on them more effectively and achieve better response performance. However, when the positive units are randomly scattered throughout the knowledge base, LLMs often experience the “Lost in the Middle” Liu et al. (2024) phenomenon, making them more likely to respond incorrectly. This suggests that for structured knowledge retrieval scenarios, recalling positive units as early as possible can effectively enable LLMs to focus on them, thereby improving performance.
In Fig. 7 and 8, we present the mean and standard deviation of F1 scores, and their correlation matrix across different subsets under the order insensitivity testbed. As illustrated in Fig. 8, we can observe that the order sensitivity of LLMs across 4 SK types exhibits a positive correlation, and so does their F1 performance. From the perspective of standard deviation, models that are insensitive to the order of SK are generally either weaker-performing LLMs, such as Llama3.1-8B, or exceptionally strong-performing LLMs, such as DeepSeek-R1. The former consistently exhibits weaker capabilities across various order settings, while the latter demonstrates stronger understanding and reasoning abilities, suggesting that current LLMs are evolving towards greater robustness and less sensitive to the order of knowledge units.
Model KG Table KG+Text Table+Text Avg.
Open Source LLMs
Llama3.1-8B 49.36 47.46 48.21 56.57 50.40
TableGPT-2 83.69 70.85 85.13 93.94 83.40
Qwen2.5-7B 81.55 70.17 75.90 80.81 77.11
GLM4-9B 69.96 61.69 63.59 71.72 66.74
Mistral-7B 61.37 62.71 51.28 53.03 57.10
Advanced General-Purpose LLMs
DeepSeek-v3 78.54 69.83 58.97 69.70 69.26
GPT-4o 87.98 73.56 76.92 80.81 79.82
DeepSeek-R1 91.42 72.88 68.21 82.32 78.71
Table 5: Negative Rejection results on four subsets.
4.4 Information Integration Analysis
From the results shown in Fig. 9, it can be observed that as the number of knowledge units required increases, the overall performance of the LLMs tends to decline. This phenomenon is more pronounced in the KG and KG+Text subsets. We believe this is due to the fact that noisy knowledge units and positive knowledge units in the KG are derived from subgraph. Many noisy units share the same entities or relations as the positive units and exhibit higher semantic similarity, which more significantly impacts the LLM’s understanding. In contrast, the row units of table data are relatively more semantically independent, so this downward trend is less noticeable in the Table subset.
In terms of understanding heterogeneous data, it is evident that as the volume of heterogeneous data increases, the performance of most LLMs declines quite substantially. Notably, in the Table+Text subset with >4 heterogeneous units, the advanced LLMs such as DeepSeek-R1 and GPT-4o still maintain relatively strong performance, whereas smaller LLMs like TableGPT-2 and Llama3.1-8B struggle to generate correct answers. Thus, we consider enhancing the ability of smaller LLMs to understand heterogeneous data to be a promising research direction worthy of further exploration.
4.5 Negative Rejection Analysis
The results in Table 5 present the rejection rates when only noisy knowledge units are provided. Overall, there is a certain positive correlation between the structured knowledge understanding performance of the LLMs and its negative rejection ability. However, we find that even DeepSeek-R1, with a negative rejection rate of 78.71%, remains vulnerable to noise interference. To our surprise, compared to Qwen2.5-7B, TableGPT-2 after fine-tuning with table-specific instructions, demonstrates stronger negative rejection ability, even surpassing GPT-4o and DeepSeek-R1. Therefore, how to strike a balance between improving the LLM’s performance and enhancing its negative rejection ability remains challenging.
5 Conclusion
In this paper, we introduce a fine-grained structured knowledge (SK) understanding benchmark, SKA-Bench, designed to provide a more comprehensive and rigorous evaluation for LLMs in understanding SK. The instances in SKA-Bench consist of a question, an answer, positive knowledge units, and noisy knowledge units, offering greater flexibility and scalability. Through varying the order and scale of knowledge units within the knowledge base, we construct four specialized testbeds to evaluate key capabilities: Noise Robustness, Order Insensitivity, Information Integration, and Negative Rejection. Empirical results demonstrate that even powerful LLMs like GPT-4o and DeepSeek-R1 still lack comprehensive understanding and reasoning capabilities for SK. Their performance is significantly influenced by factors such as the amount of noise, order of knowledge units, and hallucinations.
Limitations
Although SKA-Bench offers a more comprehensive and rigorous benchmark for evaluating structured knowledge understanding of LLMs, certain limitations warrant careful consideration, as summarized below. (1) SKA-Bench is limited to English only and does not yet capture the performances of LLMs in understanding structured knowledge across multiple languages. (2) Constrained by resource limitations, although our SKA-Bench instances have the capability to construct longer structured knowledge bases (even >64k tokens), we have not yet explored the performance of LLMs at this scale.
Ethics Statement
In this paper, we construct SKA-Bench, which is expanded and modified based on the existing 6 structured knowledge understanding evaluation datasets. Moreover, we incorporate manual annotation and manual synthetic data verification to ensure that it does not violate any ethics.
Appendix A Original Datasets Details
We provide a brief description of all the original structured knowledge understanding datasets we used and licenses below:
  • WebQSP Yih et al. (2016). WebQuestionsSP (WebQSP) is a semantic parse-based KBQA dataset with 4,737 questions coupled with SPARQL queries for KB question answering. The answers can be extracted through executing SPARQL queries on Freebase. The dataset is released under the Microsoft Research Data License Agreement.
  • CWQ Talmor and Berant (2018). ComplexWebQuestions (CWQ) is created on top of WebQSP dataset with the intention of generating more complex (by incorporating compositions, conjunctions, superlatives or comparatives) questions in natural language. It consists of 34,689 examples, divided into 27,734 train, 3,480 dev, 3,475 test. And test set in original CWQ dataset does not contain “answer”. The whole software is licensed under the full GPL v2+.
  • WTQ Pasupat and Liang (2015). WikiTableQuestions (WTQ) is a widely used table question answering (TableQA) dataset of 22,033 complex questions with average 2.14 hop on Wikipedia tables. The dataset is released under the Apache-2.0 license.
  • TableBench Wu et al. (2025). TableBench is a comprehensive and complex benchmark, including 886 samples in 18 fields within four major categories of TableQA capabilities. The tables in TableBench have an average of 6.68 columns and 16.71 rows, and the average reasoning steps of questions is 6.26. The dataset is released under the Apache-2.0 license.
  • STaRK Wu et al. (2024). STaRK is a large-scale semi-structure retrieval benchmark on textual and relational knowledge bases, covering three domains. It consists of 263 human-generated questions and 33,627 synthesized questions. And this dataset is released under the MIT license.
  • HybridQA Chen et al. (2020b). HybridQA is a question answering dataset based on heterogeneous knowledge, and each question is aligned with a Wikipedia table and multiple free-form corpora linked with the entities in the table. The questions are collected from crowd-workers, and designed to aggregate both table and text information, which means the lack of either form would render the question unanswerable. The dataset is released under the MIT license.
Appendix B Dataset Construction Details
The annotation guideline for “Iterative Positive Units Annotation” is shown in the Fig. 10.
Moreover, we have presented specific examples of the part where LLMs are involved in the entire dataset construction process. Check satisfaction by LLM in Iterative Positive Units Annotation stage is shown in Fig. 11. Noisy Synthesis process is shown in Fig. 12. And “condition” of positive knowledge units summarizing and check satisfaction by LLMs in Noisy Units Construction stage are shown in Fig. 13 and Fig. 14.
Figure 10: The annotation guidelines for annotators.
Figure 11: The prompt for checking Positive Units.
Figure 12: The prompt for Noisy Units synthesis.
Figure 13: The prompt for “contidition” summarizing.
Figure 14: The prompt for checking Noisy Units.
Appendix C Evaluation Prompt Template
Fig. 15161718 show QA prompt templates for four subsets in Noise Robustness testbed, Order Insensitivity testbed, and Information Integration testbed. Fig. 19202122 show prompt templates of negative rejection testbed for four subsets.
Figure 15: The prompt for KG subset in QA task.
Figure 16: The prompt for Table subset in QA task.
Figure 17: The prompt for KG+Text subset in QA task.
Figure 18: The prompt for Table+Text subset in QA task.
Figure 19: The prompt for KG subset in “negative rejection” testbed.
Figure 20: The prompt for Table subset in “negative rejection” testbed.
Figure 21: The prompt for KG+Text subset in “negative rejection” testbed.
Figure 22: The prompt for KG+Text subset in “negative rejection” testbed.

### v2

@@
 Empirical evaluations on 8 representative LLMs, including the advanced DeepSeek-R1, indicate that existing LLMs still face significant challenges in understanding structured knowledge, and their performance is influenced by factors such as the amount of noise, the order of knowledge units, and hallucination phenomenon.
-Our dataset and code are available at https://github.com/Lza12a/SKA-Bench.
+Our dataset and code are available at https://github.com/zjukg/SKA-Bench.
 SKA-Bench: A Fine-Grained Benchmark for Evaluating Structured Knowledge Understanding of LLMs
@@
 With the rapid development of large language mo-
-dels (LLMs) OpenAI (2023); Dubey et al.
-(2024), Structured Knowledge (SK), such as knowledge graphs (KG) Bollacker et al.
+dels (LLMs) OpenAI (2023); AI@Meta (2024), Structured Knowledge (SK), such as knowledge graphs (KG) Bollacker et al.
 (2008) and tables, still remain essential due to their systematic and rigorous organizational formats.
@@
 (2023), LLMs still face significant challenges Fang et al.
+(2024); Liu et al.
 (2025) in understanding structured knowledge.
@@
 Furthermore, the annotation process need to adhere to the following requirements: (1) if the answer is wrong, delete the sample directly; (2) if the question involves multiple answers, all positive units require to obtain the answers should be annotated; (3) for the Table subset and Table+Text subset, if the question needs to perform numerical analysis on the entire table, the corresponding SKA-QA pairs should either be removed or the question should be modified; (4) if the tables in the Table subset and Table+Text subset are order-dependent (i.e., modifying the row order would result in semantic errors in the table), this sample should be removed; (5) for the KG+Text subset and Table+Text subset, if question only utilizes one type of knowledge source, the question should be modified or removed.
-After each round of annotation, we query the LLM (utilizing DeepSeek-v3 DeepSeek-AI et al.
-(2024)) to determine whether annotated positive units can derive the answer to the given question.
+After each round of annotation, we query the LLM (utilizing DeepSeek-v3 DeepSeek-AI (2024)) to determine whether annotated positive units can derive the answer to the given question.
 If the response is “No”, re-annotation is performed.
@@
 First, we leverage LLMs with existing SKA-QA instances to generate noisy units.
-To ensure the diversity of synthesized units, we alternately use GPT-4o OpenAI (2023) and DeepSeek-v3 DeepSeek-AI et al.
-(2024) during this process.
+To ensure the diversity of synthesized units, we alternately use GPT-4o OpenAI (2023) and DeepSeek-v3 DeepSeek-AI (2024) during this process.
 Meanwhile, we also need to ensure that the synthesized noisy units do not affect the correctness of the answers.
@@
 (2024) in text understanding evaluation, we construct the four testbeds based on SKA-Bench instances to evaluate the following fundamental capabilities of LLMs in structured knowledge (SK) understanding:
-[…]
+$\bullet$ Noise Robustness.
+Here, we define noise as the remaining triples in the KG subgraph or the irrelevant rows in the table.
+We incorporate noise units of varying proportions into the positive knowledge units as the knowledge base to evaluate whether the LLM can robustly provide accurate answers.
+Considering the differences in the token counts across different knowledge units, we use the total token length as the split standard to construct test sets.
+Specifically, we construct four test sets {1k, 4k, 12k, 24k} for the Table and KG subsets, and three test sets {4k, 12k, 24k} for the Table+Text and KG+Text subsets, with the detailed statistics shown in Table 2.
+Additionally, to eliminate the influence of the knowledge unit order, we randomly shuffle the SK units in the KG and text units with a random seed of 42, while preserving the original order of the SK units in the Table.
 Subset #num SK #num unSK #token
@@
 Cells with darker colors indicate the better performance under this subset.
-[…]
-[…]
+$\bullet$ Order Insensitivity.
+SK representation naturally does not depend on any specific order.
+And in retrieval-augmented scenarios Fan et al.
+(2024), the order of retrieved knowledge units tends to be disrupted.
+Therefore, we expect LLMs to be order-insensitive when understanding SK and capturing the semantic relationships between SK units.
+In this testbed, we provide SK bases with different permutations of SK units to test whether the LLM is sensitive to order.
+For SK units in KG and textual units, we position the positive knowledge units at the beginning, randomized positions, and the end of the knowledge base, denoting them as {prefix, random, suffix}.
+For SK units in Table, we additionally introduce the original table order, denoted as {original, prefix, random, suffix}.
+Furthermore, we standardize the test sets to a scale of 4k tokens for Table and KG subsets, and 12k for Table+Text and KG+Text subsets.
+$\bullet$ Information Integration.
+This ability requires LLMs to integrate multiple knowledge units to answer questions, including the integration of multiple SK units and the integration of heterogeneous data (SK+Text) units.
+Therefore, this testbed focuses on analyzing the performance of LLMs under these two settings.
+Specifically, we divide our dataset based on the number of knowledge units required to answer each question {2, 3, 4, more than 4} to evaluate the information integration capability of LLMs.
+Regarding dataset scale and order, we standardize the test set to a scale of 4k tokens for the Table and KG subsets, and 12k tokens for Table+Text and KG+Text subsets.
+Meanwhile, we randomly shuffle (with random seed 42) the SK units in KG and text units while preserving the original order of SK units in the Table subset.
 Figure 5: Overall noise robustness results on four subsets. ‘Average’ represents the average results across on all results of four subsets.
@@
 Our evaluation is based on popular large language models (LLMs) with a context window of at least 24k tokens.
-Our evaluated LLMs include advanced general-purpose LLMs: DeepSeek-v3 DeepSeek-AI et al.
-(2024), GPT-4o OpenAI (2023), DeepSeek-R1 DeepSeek-AI et al.
-(2025) and common open-source LLMs: Llama-3.1-8B-Instruct Dubey et al.
-(2024), Qwen2.5-7B-Instruct Team (2024), GLM4-9B-Chat Zeng et al.
+Our evaluated LLMs include advanced general-purpose LLMs: DeepSeek-v3 DeepSeek-AI (2024), GPT-4o OpenAI (2023), DeepSeek-R1 DeepSeek-AI (2025) and common open-source LLMs: Llama-3.1-8B-Instruct AI@Meta (2024), Qwen2.5-7B-Instruct Team (2024), GLM4-9B-Chat Zeng et al.
 (2024), and Mistral-7B-Instruct-v0.3 Jiang et al.
@@
 (2) Constrained by resource limitations, although our SKA-Bench instances have the capability to construct longer structured knowledge bases (even >64k tokens), we have not yet explored the performance of LLMs at this scale.
+(3) Potential biases introduced by dataset selection are inevitable.
+We have made effort to minimize this impact through careful dataset selection and evaluation design.
+For dataset selection, SKA-Bench adopts widely used and well-recognized benchmark datasets for each respective data type (e.g., WebQSP, CWQ, WTQ, TableBench, StarK-Amazon, and HybridQA).
+These datasets focus on general domains and do not require extensive domain-specific knowledge, which helps to mitigate the bias introduced by domain dependence.
+For evaluation design, our primary focus is on whether LLMs can comprehensively understand various types of structured knowledge, rather than investigating the “performance differences between different modalities”.
+(4) Beyond its role as a structured knowledge understanding dataset, SKA-Bench can be effectively extended to various task scenarios, such as Text2Query (i.e., Text2SQL Guo et al.
+(2019) and Text2SPARQL Hu et al.
+(2018)), structured knowledge retrieval Mahalingam et al.
+(2024), and evaluations of knowledge-augmented systems Liang et al.
+(2024); Wang et al.
+(2025).
+These potential evaluation directions remain unexplored.
 Ethics Statement
 In this paper, we construct SKA-Bench, which is expanded and modified based on the existing 6 structured knowledge understanding evaluation datasets.
-Moreover, we incorporate manual annotation and manual synthetic data verification to ensure that it does not violate any ethics.
+These datasets have stated that there are no ethical concerns.
+Moreover, we also incorporate manual annotation and manual synthetic data verification to ensure that it does not violate any ethics.
+Acknowledgments
+This work is founded by National Natural Science Foundation of China (NSFC62306276/NSFCU23B2055/NSFCU19B2027), Zhejiang Provincial Natural Science Foundation of China (No.
+LQ23F020017), Yongjiang Talent Introduction Programme (2022A-238-G), and Fundamental Research Funds for the Central Universities (226-2023-00138).
+This work was supported by Ant Group.
 Appendix A Original Datasets Details
@@
 The dataset is released under the MIT license.
-Appendix B Dataset Construction Details
+Type 1-hop 2-hop 3-hop $\mathbf{\geq}$4-hop
+KG subset 7.59% 22.07% 16.26% 54.09%
+KG+Text subset 7.04% 10.91% 9.76% 72.29%
+Table 6: Hop count distributions in KG subset and KG+Text subset.
+Model KG KG+Text
+Hard Noise Mixed Noise Easy Noise Std. Hard Noise Mixed Noise Easy Noise Std.
+LLaMA3.1-8B 55.12 58.19 60.23 2.10 65.68 67.02 66.96 0.62
+Qwen2.5-7B 58.36 60.00 67.82 4.13 74.96 76.51 77.15 0.92
+DeepSeek-v3 71.09 73.93 74.28 1.43 77.01 77.12 77.65 0.28
+GPT-4o 72.10 73.42 74.82 1.11 76.58 77.38 77.63 0.45
+Table 7: Performance of four representative LLMs under different noise intensities on KG and KG+Text subsets.
+Appendix B Effects of Noise Distribution in KG-related Subsets
+Due to the structural complexity of knowledge graphs, the degree of noise can vary across different triples.
+We hypothesize that hop count is closely related to semantic similarity, and this relationship may affect the noise distribution, thereby impacting model performance.
+To investigate this, we conduct an in-depth analysis of the hop count distribution in the existing KG subsets and KG+Text subsets, and provide experimental results to support this hypothesis.
+As shown in Table 6, we present the detailed hop count distributions for the KG and KG+Text subsets.
+It is worth noting that we report the actual distributions in SKA-Bench rather than the original datasets, since some subgraphs were randomly truncated during the graph extraction process.
+Thanks to the high flexibility of SKA-Bench, we further present results under different noise intensities to more thoroughly evaluate the noise robustness of LLMs.
+In our experimental setup, the context size is consistently controlled at 4k, and knowledge units from the knowledge graph are arranged in a random order.
+Noise is categorized as “Hard Noise” ($\leq$2 hops), “Mixed Noise” (same as in main text), and “Easy Noise” ($\geq$4 hops).
+We report the performance of four representative LLMs, as shown in Table 7.
+The experimental results show that the variation in model performance is significantly correlated with the intensity of noise.
+Comparatively, stronger LLMs demonstrate greater robustness to noise, and this conclusion also holds when the noise level is expanded horizontally (as in the noise robustness testbed of the main text).
+Appendix C Dataset Construction Details
 The annotation guideline for “Iterative Positive Units Annotation” is shown in the Fig. 10.
@@
 Figure 14: The prompt for checking Noisy Units.
-Appendix C Evaluation Prompt Template
+Appendix D Evaluation Prompt Template
 Fig.